The decay rate of the carbon in an old bone is one-eighth that of the carbon in a living organism. The half-life of carbon-14 is 5730 yr. How old is the bone
1
Expert's answer
2018-03-14T08:48:04-0400
In half-life period activity decay in two time. In several periods half-life it decay in (1/2)^n, where n-number periods half-life.

From
(1/2)^n=1/8,
n=3
Bone is 3*5730=17190 year old